OpenStudy (anonymous):

Determine the solution to the system of equations given below. y=x^2-6x+10 y=-x+4 A.(2,3) B.(3,1) C.(2,2) and (3,1) D. no solution

OpenStudy (anonymous):

use substitution :)

OpenStudy (satan ):

Try solving them: \[\Large -x+4=x^2-6x+10\]

OpenStudy (satan ):

Get it equal to zero: \[\Large 0=x^2+5x+6\] Factoring anyone?

OpenStudy (satan ):

Whoops, *-5x not +5x

OpenStudy (anonymous):

Is x=-1

OpenStudy (anonymous):

it will turn into a quadratic equation

OpenStudy (anonymous):

\[x ^{2}-5x +6=0\]

OpenStudy (anonymous):

I got (3,2)

OpenStudy (anonymous):

c